For an electron in the l = 2 state:

(a) enumerate all the possible values of quantum numbers j and mj ;

(b) draw the corresponding vector diagrams;

(c) estimate the maximum value of the spin-orbit coupling energy E. 


1
Expert's answer
2021-05-20T10:09:31-0400

"l=2\\\therefore n=3"

(a) "s=1\/2"

"j=(n+s),(n-s)"

"j=(5\/2),(3\/2)"


For j = 5/2

"jm=j(j+1)\hbar^2"

"jm=\dfrac{35}{4}\hbar^2"


For j = 3/2

"jm=\dfrac{15}{4}\hbar^2"


(c) "\Delta E_{so}=\alpha\dfrac{Z\hbar^3}{4m^2c}\dfrac{1}{r^3}[j(j+1)-l(l+1)-s(s+1)]"

"s=1/2,\space l=2,\space j=5/2"

"\Delta E_{so}=\alpha\dfrac{Z\hbar^3}{4m^2c}\dfrac{1}{r^3}\bigg[\dfrac{5}{2}\bigg(\dfrac{5}{2}+1\bigg)-2(2+1)-\dfrac{1}{2}\bigg(\dfrac{1}{2}+1\bigg)\bigg]"

"\Delta E_{so}=2\alpha\dfrac{Z\hbar^3}{4m^2c}\dfrac{1}{r^3}"